OpenStudy (anonymous):

using the point-slope equation, find the equation containing (-5,-8) and slope m= -3

OpenStudy (lgbasallote):

use the point slope form \[y - y_1 = m(x-x_1)\] m = 3 x1 = -5 y1 = -8